the perimeter of a rectangle is 36 yd. the width is 18 yds less han twice the lenght. find the length and the width of the rectangle
Answer by checkley79(3341)

P=2L+2W
W=2L-18
36=2L+2(2L-18)
36=2L+4L-36
36=6L-36
6L=36+36
6L=72
L=72/6
L=12 ANS. FOR THE LENGTH.
W=2*12-18
W=24-18
W=6 ANS. FOR THE WIDTH.
PROOF:
36=2*12+2*6
36=24+12
36=36